At 0.31 per bushel, the daily supply for wheat is 378 bushels, and the daily demand is 514 bushels. When the price is raised to $0.79 per bushel the daily supply increases to 538 bushels, and the daily demand decreases to 434 bushels. Assume that the price-supply and price-demand equations are linear
Answer by Boreal(15235)

Write as ordered pairs, using cents for x: (31,378)(79,538)
Slope is (160/48)=10/3
use point slope formula y-y1=m(x-x1), m slope, (x1,y1) one of the points.
y-378=(10/3)(x-31)
y-378=(10x/3)-(310/3); 310/3=103 1/3
y=(10x/3)+274 2/3: supply equation
-----------------------
(31,514)(79,434)
slope is -80/48=-5/3
y-514=(-5/3)(x-31)
y-514=(-5x/3)+155/3
y=(-5x/3)+565 2/3